# Modeling local predictive ability using power-transformed Gaussian processes

### v1

@@ 2.4 Cube root transformed Gaussian process regression
HMC sampling for the scaled non-central $\chi^{2}$ Gaussian process regression is efficient, but necessarily slow in even moderately large datasets since $\boldsymbol{\lambda}_{k}$ is high-dimensional. This subsection presents a cube root transformation of the log scores which makes the log scores approximately normal with constant variance. This allows us to analytically integrate out the high-dimensional latent ELPD surfaces from the posterior, thereby substantially speeding up the computation of the ELPD posterior at a new point in the pooling space. We will refer to this model in tables and figures as $\operatorname{GP}\left(\nicefrac{{1}}{{3}}\right)$.
Starting from the linear transformation $\ell^{\prime}_{ik}=-(\ell_{ik}-a_{ik})$, we further transform the log scores $\ell_{ik}^{\prime}$ by taking the cube root, which is motivated by Abdel-Aty, (1954) who shows that a version of the Wilson-Hilferty transformation of the $\chi^{2}_{k}$ distribution approximately holds for $X\sim\chi_{k}^{2}(\lambda)$ as well. That is
$\left(\frac{X}{k+\lambda}\right)^{1/3}\dot{\sim}\,\mathcal{N},$ (10)
where the mean and variance of the normal distribution depends on the non-centrality parameter $\lambda$ as well as the degrees of freedom. Since the non-centrality parameter $\lambda$ is not known, we use a plain cube-root $X^{1/3}$ transformation when transforming the log scores. Figure 2 shows that the this transformation makes the log scores approximately Gaussian; compare with Figure 1. As $\lambda$ goes towards infinity the variance goes to zero, but Figure 3 shows that the variance is approximately constant in a range of reasonable $\lambda$ values. We use the notation $\ell^{\prime\prime}$ for these twice transformed log scores.
@@
$\eta_{k}(\tilde{\mathbf{z}})=\tilde{a}_{k}-f_{k}^{3}(\tilde{\mathbf{z}})-3f_{k}(\tilde{\mathbf{z}})\sigma_{k}^{2}$ (14)
@@
where $\boldsymbol{\ell}_{k}^{\prime\prime}$ is a vector with transformed log scores in the training data. This low-dimensional marginal posterior can be sampled with Hamiltonian Monte Carlo (HMC) using Stan (Stan Development Team,, 2022; Gabry and Češnovar,, 2020) in a fraction of the time that it takes to sample with full joint posterior for the $\mathrm{GP}(\chi_{1}^{2})$ model in (9).
@@
Input: Number of posterior draws $M$.
@@
       Compute the ELPD draw: $\eta_{k}^{(j)}(\tilde{\mathbf{z}})=a_{*k}-(f_{k}^{(j)}(\tilde{\mathbf{z}}))^{3}-3f_{k}^{(j)}(\tilde{\mathbf{z}})(\sigma_{k}^{(j)})^{2}$
@@
Output: Sample from ELPD posterior: $\eta_{k}^{(1)},\eta_{k}^{(2)},\ldots,\eta_{k}^{(M)}$.
Algorithm 1 Sampling from the ELPD posterior in the $\operatorname{GP}\left(\nicefrac{{1}}{{3}}\right)$ model
@@ 2.5 General power-transformed Gaussian process regression
The cube root transformation is suitable when the log scores follow a scaled non-central $\chi^{2}_{1}$ distribution, which is the correct model in the important case when both the DGP and the expert is Gaussian. Deviations from Gaussianity lead to new models that need to be derived on a case by case basis. The posterior distribution of the ELPD can be sampled with HMC as in the non-central $\chi^{2}_{1}$ model described above. There may be new transformations that convert the log scores to something that can be modeled by a Gaussian process regression with Gaussian errors, but also this needs to be determined for each particular model.
@@
3 Simulation study
To explore the properties of our methods we implement them on simulated data. We are particularly interested in investigating how the computationally faster $\operatorname{GP}(\nicefrac{{1}}{{3}})$ model compares to the computationally costly $\operatorname{GP}(\chi_{1}^{2})$ model.
We generate data from the model
We further assume that our decision maker believes that the predictive ability of the expert varies over the variable $x_{2}$, but does not know the exact relationship between the ELDP $\eta$ and $x_{2}$. We will examine how well the decision maker will be able to estimate the true predictive ability, as summarized by the posterior distribution of $\eta$, using the two models described in Section 2.
The true distribution of the log scores in this setup is given by
@@ 3 Simulation study
Refer to caption Refer to caption
Figure 5: The left figure shows the difference in log mean integrated squared error (MISE) of the $\operatorname{GP}\left(\nicefrac{{1}}{{3}}\right)$ and $\operatorname{GP}(\chi^{2}_{1})$ model, taken over the distribution of $x_{2}$.
The right figure shows the difference in mean integrated log score (MILS) of the $\operatorname{GP}(\nicefrac{{1}}{{3}})$ and $\operatorname{GP}(\chi^{2}_{1})$ model, taken over the distribution of $x_{2}$.
To compare the $\operatorname{GP}\left(\chi^{2}_{1}\right)$ and $\operatorname{GP}\left(\nicefrac{{1}}{{3}}\right)$ models we simulate $1000$ data sets of $150$ observations each from the DGP in Equation (17). For each simulation, we then generate draws from the posterior distribution of $\eta(x_{2})$ over a grid of $x_{2}$ values between $-3$ and $3$, and compare with the true ELPD as a function of $x_{2}$. Both GPs use a squared exponential kernel (Rasmussen and Williams,, 2006). We use $l\sim\operatorname{Inverse-Gamma}(5,5)$ priors for the length scales, and a $\alpha\sim\mathcal{N}^{+}(0,1)$ prior for the signal variances. The noise variance $\sigma_{n}^{2}$ of the $\operatorname{GP}(\nicefrac{{1}}{{3}})$ is given a $\mathcal{N}^{+}(0,1)$ prior as well.
Figure 4 show the posterior mean of $\eta(x_{2})$ together with a $95\%$ credible interval for three different data sets. The data sets have been selected by ranking the mean integrated squared error (MISE) of the ELPD estimates of each run – the MSE integrated over the distribution of $x_{2}$ – and then selecting the $2.5$:th, $50$:th, and $97.5$:th percentile. Both methods perform well for $x_{2}\in(-2,2)$, but poorly in the tails. This is to be expected, as two factors conspire to make estimation in the tails difficult: the DGP variance of the log scores increases with the square of $x_{2}$, and since $x_{2}\sim\mathcal{N}(0,1)$ only roughly $5\%$ of the observations are in the tails.
Note that the uncertainty around the estimate increases quickly for both models when there are no, or few, previous predictions nearby. This will lead to the positive behavior that when combining experts in regions where we have little to go on in terms of previous predictions, we will quickly tend towards an equal-weights solution; see Section 4.
The $\operatorname{GP}(\chi^{2}_{1})$ is prohibitively costly from a computational standpoint once the number of observations goes above a few hundred. The fact that the more scalable $\operatorname{GP}\left(\nicefrac{{1}}{{3}}\right)$ model works so well compared to the $\operatorname{GP}(\chi^{2}_{1})$ model is therefore an important finding for practical applications. There is also a large collection of methods for reducing computational cost in GP regression with additive homoscedastic errors (Liu et al.,, 2020) that can be directly used to speed up inference in the $\operatorname{GP}\left(\nicefrac{{1}}{{3}}\right)$ model.
4 Application to bike rental prediction
Rental services for bikes or similar vehicles is a large and growing business in many developed urban areas. An important task for such businesses is the prediction of vehicles usage, on a daily or more granular basis. This section applies the method described in Section 2 to estimate the local one-step-ahead predictive ability of three experts/models predicting bike rentals based on the data in Fanaee-T and Gama, (2014). We then propose and implement a method for aggregating predictive distributions based on these estimates.
4.1 Inferring local predictive ability

### v2

@@ 2.4 Cube root transformed Gaussian process regression
HMC sampling for the scaled noncentral $\chi^{2}$ Gaussian process regression is efficient, but necessarily slow in even moderately large datasets since $\boldsymbol{\lambda}_{k}$ is high-dimensional. This subsection presents a cube root transformation of the log scores which makes the log scores approximately normal with constant variance. This allows us to analytically integrate out the high-dimensional latent ELPD surfaces from the posterior, thereby substantially speeding up the computation of the ELPD posterior at a new point in the pooling space. We will refer to this model in tables and figures as $\operatorname{GP}\left(\nicefrac{{1}}{{3}}\right)$.
Starting from the linear transformation $\ell^{\prime}_{ik}=-(\ell_{ik}-a_{ik})$, we further transform the log scores $\ell_{ik}^{\prime}$ by taking the cube root, which is motivated by Abdel-Aty, (1954) who shows that a version of the Wilson-Hilferty transformation of the $\chi^{2}_{v}$ distribution approximately holds for $X\sim\chi_{v}^{2}(\lambda)$ as well. That is
$\left(\frac{X}{v+\lambda}\right)^{1/3}\dot{\sim}\,\mathcal{N},$ (10)
where the mean and variance of the normal distribution depends on the noncentrality parameter $\lambda$ as well as the degrees of freedom. Since the noncentrality parameter $\lambda$ is not known, we use a plain cube-root $X^{1/3}$ transformation when transforming the log scores. Figure 2 shows that the this transformation makes the log scores approximately Gaussian; compare with Fig. 1. As $\lambda$ goes towards infinity the variance goes to zero, but Fig. 3 shows that the variance is approximately constant in a range of reasonable $\lambda$ values. We use the notation ′′$\ell^{\prime\prime}=(\ell^{\prime})^{1/3}$ for these twice transformed log scores.
@@
$\eta_{k}(\tilde{\mathbf{z}})=\tilde{a}_{k}-f_{k}^{3}(\tilde{\mathbf{z}})-3f_{k}(\tilde{\mathbf{z}})\sigma_{k}^{2}\;.$ (14)
@@
where $\boldsymbol{\ell}_{k}^{\prime\prime}$ is a vector with transformed log scores in the training data. This low-dimensional marginal posterior can be sampled with Hamiltonian Monte Carlo (HMC) using Stan (Stan Development Team,, 2024; Gabry and Cesnovar,, 2022) in a fraction of the time that it takes to sample with full joint posterior for the $\mathrm{GP}(\chi_{1}^{2})$ model in (9).
@@
Input: Number of posterior draws $M$, vector of pooling variables $\tilde{\mathbf{z}}$ and $\tilde{a}_{k}=-(1/2)\log(2\pi\tilde{\sigma}_{k}^{2})$ for expert $k$.
@@
       Compute the ELPD draw: $\eta_{k}^{(j)}(\tilde{\mathbf{z}})=\tilde{a}_{k}-(f_{k}^{(j)}(\tilde{\mathbf{z}}))^{3}-3f_{k}^{(j)}(\tilde{\mathbf{z}})(\sigma_{k}^{(j)})^{2}$
@@
Output: Sample from ELPD posterior: $\eta_{k}^{(1)}(\tilde{\mathbf{z}}),\eta_{k}^{(2)}(\tilde{\mathbf{z}}),\ldots,\eta_{k}^{(M)}(\tilde{\mathbf{z}})\;.$
Algorithm 1 Posterior sampling for the ELPD at $\tilde{\mathbf{z}}$ in the $\operatorname{GP}\left(\nicefrac{{1}}{{3}}\right)$ model
As the $\operatorname{GP}(\nicefrac{{1}}{{3}})$ adds an approximation layer on top of the original model, we want to ensure that the gains in computational speed do not come at the cost of a drop in performance. To this end, we conduct a simulation study that shows good performance of the $\operatorname{GP}(\nicefrac{{1}}{{3}})$ model compared to the $\operatorname{GP}(\chi^{2}_{1})$ model. For details, see Appendix.
@@ 2.5 General power-transformed Gaussian process regression
The cube root transformation is suitable when the log scores follow a scaled noncentral $\chi^{2}_{1}$ distribution, which is the correct model in the important case when both the DGP and the expert is Gaussian. Deviations from Gaussianity lead to new models that need to be derived on a case by case basis. The posterior distribution of the ELPD can be sampled with HMC as in the noncentral $\chi^{2}_{1}$ model described above. There may be new transformations that convert the log scores to something that can be modeled by a Gaussian process regression with Gaussian errors, but also this needs to be determined for each particular model.
@@
2.6 The multi-output $\operatorname{GP}(\nicefrac{{1}}{{3}})$ model
The $\operatorname{GP}(\nicefrac{{1}}{{3}})$ model proposed in Sect. 2.4 can be used to model the predictive ability of each of a set of experts. If these estimates are then used to create a pooled predictive distribution two independence assumptions are implied: independence in the predictive abilities $\eta_{1}(\mathbf{z}_{i}),\ldots,\eta_{K}(\mathbf{z}_{i})$, and independence in the errors, $\ell^{\prime\prime}_{ik}-f_{k}(\mathbf{z}_{i})$ for $i=1,\ldots,n$, conditional on the predictive ability. In this section, we propose a generalisation of the $\operatorname{GP}(\nicefrac{{1}}{{3}})$ model to a multi-output GP that relaxes both of these assumptions.
Let $\ell_{ik}^{\prime\prime}=\left(a_{ik}-\ell_{ik}\right)^{\nicefrac{{1}}{{3}}}$ denote the cube-root transformed log scores. The distribution of $\ell_{ik}^{\prime\prime}$ for $K$ experts jointly, $\boldsymbol{\ell}_{i}^{\prime\prime}=(\ell_{i1}^{\prime\prime},\ldots,\ell_{iK}^{\prime\prime})^{\top}$, can then be modeled using a multi-output GP with Gaussian noise
$\begin{split}\boldsymbol{\ell}^{\prime\prime}_{i}\mid\mathbf{f}(\mathbf{z}_{i}),\boldsymbol{\Sigma}&\sim\mathcal{N}\left(\mathbf{f}(\mathbf{z}_{i}),\boldsymbol{\Sigma}\right)\\ \mathbf{f}(\mathbf{z})\mid\boldsymbol{\Theta},\mathbf{C}&\sim\operatorname{GP}\left(\boldsymbol{\mu}(\mathbf{z}),g(\mathbf{z},\mathbf{z}^{\prime})\right),\end{split}$ (17)
where $\mathbf{f}(\mathbf{z})=(f_{1}(\mathbf{z}),\ldots,f_{K}(\mathbf{z}))^{\top}$ is a multi-output GP with hyperparameters $\boldsymbol{\Theta}=\{\boldsymbol{\theta}_{1}$,…,$\boldsymbol{\theta}_{K}\}$ including a vector of parameters $\boldsymbol{\theta}_{k}$ for the mean and covariance function of each expert as well, and $\mathbf{C}$ is a $K\times K$ matrix that will be used below to correlate the experts’ abilities $\mathbf{f}$. The noise covariance matrix $\boldsymbol{\Sigma}$ is here taken to be a general positive definite matrix, but may also be restricted to a diagonal matrix. The goal of the model in (17) is to capture correlation of the log scores between experts. This correlation is coming from two sources: in the noise distribution via $\boldsymbol{\Sigma}$, and in the underlying GP prior for $\mathbf{f}(\mathbf{z})$ with correlation generated by $\mathbf{C}$, which we now describe.
The model in (17) for all log scores in the training sample can be written
′′$\underset{n\times K}{\boldsymbol{\ell}^{\prime\prime}}$ $=\underset{n\times K}{\mathbf{F}}+\underset{n\times K}{\mathbf{E}},$ (18)
where $\boldsymbol{\ell}^{\prime\prime}$ is the matrix consisting of the transformed log scores of $K$ experts at $n$ time points, $\mathbf{F}$ is determined by the multi-output GP, and $\mathbf{E}$ is a matrix of Gaussian noise terms.
Reformulating the model in vectorized form, we can write the vector of errors as ${\operatorname{vec}\mathbf{E}=\boldsymbol{\Sigma}\otimes\mathbf{I}_{n}}$. By letting $\boldsymbol{\Sigma}$ be an arbitrary covariance matrix we allow correlation in the error terms across experts, but assume independence between observations.
Modeling the correlation in the underlying predictive ability is done via the GP prior. A first attempt for a model for $\mathbf{F}$ is a matrix normal
$\mathbf{F}\sim\operatorname{\mathcal{MN}}(\mathbf{M},\mathbf{G},\mathbf{\Omega}),$ (19)
i.e. that $\operatorname{vec}\mathbf{F}$ follows a multivariate normal with Kronecker-structured covariance matrix $\mathbf{\Omega}\otimes\mathbf{G}$. The matrix $\mathbf{G}$ models the correlation between rows of $\mathbf{F}$, which correspond to observations, and is therefore typically a covariance matrix generated from a kernel function, for example the squared exponential in (7). The matrix $\mathbf{\Omega}$ models the correlation between columns, in our case experts, and can most naturally be taken as a general positive definite covariance matrix. The signal variances in the squared exponential kernel, $\alpha$, would then be set to $1$ for each expert.
While the matrix normal model in (19) has some intuitive appeal and is easy to work with from a computation perspective, it assumes that the columns of $\mathbf{F}$ all follow the same multivariate normal distribution, which puts the restriction on our model that all experts have identical covariance function, i.e. the same covariance matrix $\mathbf{G}$. To be able to specify separate covariance functions for each expert we instead induce a prior for $\mathbf{F}$ by correlating a set of $K$ underlying independent GPs, $\mathbf{h}_{1},\ldots,\mathbf{h}_{K}$, each with its own kernel function (Teh et al.,, 2005)
$\mathbf{F}=\mathbf{H}\mathbf{C},\qquad\mathbf{H}=\left[\mathbf{h}_{1}\dots\mathbf{h}_{K}\right],\qquad\mathbf{h}_{k}\overset{\mathrm{indep}}{\sim}\mathcal{N}\left[\mathbf{0},\operatorname{G}_{k}(\mathbf{Z},\mathbf{Z})\right],$ (20)
where $\operatorname{G}_{k}(\mathbf{Z},\mathbf{Z}^{\prime})$ is the covariance matrix for expert $k$ generated from kernel function $g_{k}(\mathbf{z},\mathbf{z}^{\prime})$ and $\mathbf{C}$ is a $K\times K$ matrix that generates the dependence between the $\mathbf{f}_{k}$ for the experts. We set the signal standard deviation to unity in all kernels, i.e. $\alpha_{1}=\cdots=\alpha_{K}=1$ in the squared exponential kernel in (7) so that the scales of $\mathbf{F}$ is generated by the $\mathbf{C}$ matrix. While the decision maker does not need to use the same pooling variables for each expert, we let the matrix of pooling variables $\mathbf{Z}$ contain the pooling variables for all expert to simplify notation.
The model in (20) allows the decision maker to specify completely different covariance functions for the experts. Importantly, this allows the length scales to differ between the expert in the case where an ARD kernel is used, but it is also possible to use, for example, a Matérn kernel (Rasmussen and Williams,, 2006) for some of the experts and squared exponential for others. It is also straightforward to force a subset of experts to share the exact same covariance function.
The covariance function for $\mathbf{F}$ in (20) is a weighted sum of the covariance functions of the $K$ experts
$\operatorname{cov}(f_{k}(\mathbf{z}_{i}),f_{l}(\mathbf{z}_{j}))=\sum_{s=1}^{K}c_{sk}c_{sl}\mathrm{g}_{s}(\mathbf{z}_{i},\mathbf{z}_{j}),$ (21)
where $c_{kl}$ is element $(k,l)$ of the matrix $\mathbf{C}$. To fully specify the multi-output $\operatorname{GP}(\nicefrac{{1}}{{3}})$ model — which we will denote by $\operatorname{multi-GP}(\nicefrac{{1}}{{3}})$ — we need to specify the covariance function of each expert, including priors for any unknown hyperparameters, as well as a prior for the matrix $\mathbf{C}$ that describes the between-expert covariances.
Note that while we choose to select a number of underlying processes equal to the number of experts, there is nothing that prevents us from creating our multi-output process by correlating an arbitrary number of latent GPs. A common justification for using several models is that it gives a more holistic quantification of uncertainty by capturing the structural uncertainty of the parametric form of the models, rather than just the parametric uncertainty within models (Draper,, 1995). From this perspective, including several models that share the majority of structural DNA can be problematic, for example in Bayesian model averaging (Hoeting et al.,, 1999), where adding an effective ”copy” of an already existing model while using a uniform prior will essentially double the posterior model weight of that model. When modeling the joint predictive ability of a large pool of highly correlated experts it is possible that many of the elements of $\mathbf{C}$ will be close to zero, in which case restricting the number of underlying GPs could simplify the model and speed up calculations.
Inference for the multi-output $\operatorname{GP}(\nicefrac{{1}}{{3}})$ model
When modeling local predictive ability jointly, the end goal is the joint posterior of the ELPD for all experts at a new point $\tilde{\mathbf{z}}$ in the pooling space, $\boldsymbol{\eta}(\mathbf{z})=(\eta_{1}(\tilde{\mathbf{z}}),\ldots,\eta_{K}(\tilde{\mathbf{z}}))^{\top}$, i.e. the mean of $\boldsymbol{\ell}$ at $\tilde{\mathbf{z}}$. Since $\operatorname{E}(\boldsymbol{\ell}_{i})=\left(\operatorname{E}(\ell_{i1}),\dots,\operatorname{E}(\ell_{iK})\right)^{\intercal}$ and $\ell_{ik}=a_{ik}-\left(\ell_{ik}^{\prime\prime}\right)^{3}$, we can use the same formula for the third moments of Gaussians as in the univariate case and obtain the local predictive ability from model (17) pointwise as
$\eta_{k}(\tilde{\mathbf{z}})=\tilde{a}_{k}-f_{k}^{3}(\tilde{\mathbf{z}})-3f_{k}(\tilde{\mathbf{z}})\boldsymbol{\Sigma}_{k,k},$ (22)
where $\boldsymbol{\Sigma}_{k,k}$ denotes the $k$th diagonal element of the noise covariance matrix, and $f_{k}(\tilde{\mathbf{z}})$ the $k$th output of the GP at $\tilde{\mathbf{z}}$.
@@ 2.6 The multi-output $\operatorname{GP}(\nicefrac{{1}}{{3}})$ model
This lower-dimensional marginal posterior can be sampled with Hamiltonian Monte Carlo (HMC) using Stan (Stan Development Team,, 2024; Gabry and Cesnovar,, 2022) in a fraction of the time that it takes to sample the full joint posterior.
For each HMC sampled $(\boldsymbol{\Theta},\mathbf{C},\boldsymbol{\Sigma})$ from (23) we can now easily sample from the distribution of $\mathbf{f}(\tilde{\mathbf{z}})$ using (Rasmussen and Williams,, 2006)
$\mathbf{f}(\tilde{\mathbf{z}})\mid\boldsymbol{\ell}^{\prime\prime},\mathbf{Z},\tilde{\mathbf{z}},\boldsymbol{\Sigma},\boldsymbol{\Theta},\mathbf{C}\sim\mathcal{N}\left[\bar{\tilde{\mathbf{f}}},\operatorname{var}(\tilde{\mathbf{f}})\right],$ (24)
where $\mathbf{Z}$ is the matrix of pooling variables in the training data, and
$\bar{\tilde{\mathbf{f}}}$ $=\boldsymbol{\mu}(\tilde{\mathbf{z}})+\operatorname{G}(\tilde{\mathbf{z}},\mathbf{Z})\left[\operatorname{G}(\mathbf{Z},\mathbf{Z})+\boldsymbol{\Sigma}\otimes\mathbf{I}_{n}\right]^{-1}(\operatorname{vec}\boldsymbol{\ell}^{\prime\prime}-\boldsymbol{\mu}),$
$\operatorname{var}(\tilde{\mathbf{f}})$ $=\operatorname{G}(\mathbf{z},\tilde{\mathbf{z}})-\operatorname{G}(\tilde{\mathbf{z}},\mathbf{Z})\left[\operatorname{G}(\mathbf{Z},\mathbf{Z})+\boldsymbol{\Sigma}\otimes\mathbf{I}_{n}\right]^{-1}\operatorname{G}(\mathbf{Z},\tilde{\mathbf{z}}),$
where, for example, $\operatorname{G}(\mathbf{Z},\mathbf{Z})$ is the $Kn\times Kn$ matrix with covariances of the form $\operatorname{cov}(f_{k}(\mathbf{z}_{i}),f_{l}(\mathbf{z}_{j}))$ for experts $k$ and $l$ and observations $i$ and $j$ in the training data. The posterior draws of $\boldsymbol{\Sigma}$ from (23) and $\mathbf{f}(\tilde{\mathbf{z}})$ from (24) can finally be inserted into (22) to obtain draws from the posterior of the local ELPD $\boldsymbol{\eta}(\tilde{\mathbf{z}})$. The method is summarized in Algorithm 2.
Input: Number of posterior draws $M$, a vector of pooling variables $\tilde{\mathbf{z}}$ and $\tilde{a}_{k}=-(1/2)\log(2\pi\tilde{\sigma}_{k}^{2})$ for expert $k$.
Generate $M$ draws $\{\boldsymbol{\Sigma}^{(j)},\boldsymbol{\Theta}^{(j)},\mathbf{C}^{(j)}\}_{j=1}^{M}$ from $p\big{(}\boldsymbol{\Sigma},\boldsymbol{\Theta},\mathbf{C}\mid\boldsymbol{\ell}^{\prime\prime}\big{)}$ in (23) using HMC.
for $j=1,\dots,M$ do
       Sample the posterior $\mathbf{f}^{(j)}(\tilde{\mathbf{z}})\mid\boldsymbol{\ell}^{\prime\prime},\mathbf{Z},\tilde{\mathbf{z}},\boldsymbol{\Sigma}^{(j)},\boldsymbol{\Theta}^{(j)},\mathbf{C}^{(j)}$ in (24)
       Compute joint local predictive ability $\boldsymbol{\eta}^{(j)}(\tilde{\mathbf{z}})$ from (22)
end for
Output: Sample from the joint local predictive ability posterior: $\boldsymbol{\eta}^{(1)}(\tilde{\mathbf{z}}),\boldsymbol{\eta}^{(2)}(\tilde{\mathbf{z}}),\ldots,\boldsymbol{\eta}^{(M)}(\tilde{\mathbf{z}})\;.$
Algorithm 2 Posterior sampling for ELPD at $\tilde{\mathbf{z}}$ in the Multi-$\operatorname{GP}\left(\nicefrac{{1}}{{3}}\right)$ model
3 Predicting bike-sharing utilization rates in Washington D.C.
The last decade has seen an explosion in companies offering short-term rentals of final mile transportation, such as bikes or electric scooters. Key for the success of such a business is to be able to correctly predict demand. In this section we use single- and multi-output GP models to estimate the joint one-step-ahead local predictive ability of three experts predicting bike rentals based on the data in Fanaee-T and Gama, (2014). We then propose novel methods for how to use the estimates of local predictive ability to form aggregate predictions, and evaluate the single- and multi-output versions with regards to out-of-sample performance.
3.1 Data and models